OpenStudy (anonymous):

Beach Channel High School’s football team scored one field goal (3 points) and gave up two touchdowns (14 points) in the first quarter. In the third quarter, they scored another field goal and gave up a safety (2 points). They gave up a final field goal in the fourth quarter. By how much did they lose?

OpenStudy (anonymous):

so would it be -1 +1-2-3=?

OpenStudy (anonymous):

or would it be 3-14-14+3-2-3=?

OpenStudy (anonymous):

-27?

OpenStudy (anonymous):

14+2+3=19 lost by 3+3+9 score goals

OpenStudy (anonymous):

How did you figure that?

OpenStudy (anonymous):

wrong sorry 14+2+3= 19 3+3=6 not 9 sorry they should of scored 25 lost by 19 cause they gave up 2 touch downs a safety and field goal

OpenStudy (anonymous):

if wrong correct me

OpenStudy (shadowfiend):

The team scored one field goal, that's +3. They gave up two touchdowns, that's -14. They scored another field goal, another +3. They gave up a safety, that's -2. They gave up a final field goal, that's -3. So you have: 3 - 14 + 3 - 2 -3 = 3 - 17 - 2 -3 = 3 - 12 = -9 So they lost by 9 points.

OpenStudy (shadowfiend):

Er... No. 3 - 17 - 2 - 3 = 3 - 22 = -19 Heh. So they lost by 19 points, like bsanchez said :)

OpenStudy (anonymous):

lol i was re doing calculation lol

OpenStudy (anonymous):

but they gave up 2 touch down wich were 14 points

OpenStudy (anonymous):

a pieace

OpenStudy (anonymous):

i mean 3 touch downs and gave up 2

OpenStudy (anonymous):

yeah 2 touch downs a field goal and a safety they gave up 7+7+2+3=19

OpenStudy (anonymous):

oh ha ha i just realized where i went wrong

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Beach Channel High School's football team scored one field goal (3 points) and gave up two touchdowns (14 points) in the first quarter. (3 - 14) In the third quarter, they scored another field goal and gave up a safety (2 points). (3 - 14) + (3 - 2) They gave up a final field goal in the fourth quarter. (3 - 14) + (3 - 2) + (-3) By how much did they lose? (3 - 14) + (3 - 2) + (-3) = 3 - 14 + 3 - 2 - 3 = -13 They were down 13 at the end of the game.

OpenStudy (shadowfiend):

Woow. Yeah, robtobey. Apparently I forgot to do addition/subtraction. Sweetness.

OpenStudy (shadowfiend):

bsanchez: I'd go ahead and take my medal away hehe. I was completely wrong! :)

OpenStudy (anonymous):

Final score: 19 for the opposing team and 6 for Beach Channel High School.